OpenStudy (anonymous):

a 8=_3 b 4=_5 c 9=_2 d 6=_3 e 6=_5 i need help with this a.s.a.p

OpenStudy (anonymous):

what are these?

OpenStudy (anonymous):

what do u mean,. what are these????? it's my math

OpenStudy (anonymous):

what does _3 mean exactly?

OpenStudy (anonymous):

i dont know what you are asking

OpenStudy (anonymous):

Yes. can you make the question more clear @afranklin12

OpenStudy (anonymous):

yes i can. 8= somthing 3

OpenStudy (anonymous):

so do you want the"something"= ??

OpenStudy (anonymous):

8 = (something) x 3 you mean this?

OpenStudy (anonymous):

yes that is what i need

OpenStudy (anonymous):

something = 8/3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Similtarly, b= 5 / 4 c= 2 /9 d = 3/6 e= 5/6

OpenStudy (anonymous):

yes @shivam_bhalla is correct let "n" be our "something" then we have 8 = n x 3 so dividing both sides by 3 we get n = 8/3

OpenStudy (anonymous):

thanks. that was all. thanks and thanks 4 the answer